OpenStudy (anonymous):

4/5x - 3 3/4x = 2 1/2

OpenStudy (anonymous):

Do you mean \[\frac45 x-3\frac34 x=2\frac12\]

OpenStudy (anonymous):

Yeah! How do you do that?

OpenStudy (anonymous):

How i write equation? or How i solve this equatuon ?

OpenStudy (anonymous):

Both. Haha

OpenStudy (anonymous):

\[\begin{aligned} \frac45 x -3 \frac34 x &=2\frac12 \\ (\frac45 -3\frac34 )x &=2\frac12 \end{aligned}\] Can you see how to finish it?

hero (hero):

\[x = \frac{2.5}{\frac{4}{5} - 3\frac{3}{4}}\]

OpenStudy (anonymous):

No.. I really don't see how. :l

hero (hero):

I'm sorry to hear that

OpenStudy (anonymous):

2.5/ 2.95 Is that right?

OpenStudy (anonymous):

Help!!!!!

hero (hero):

What you want to end up with is either a whole number or decimal for x. It should not be in fraction form.

OpenStudy (anonymous):

Fractions are the only option choices.

hero (hero):

I see. Well, you should try to simplify the fraction and see if it matches one of your answer choices.

OpenStudy (anonymous):

Fuuuuuuhhhh! I dunno! D: what's the answer!?

hero (hero):

If you don't know how to simplify fractions, you're in big trouble

OpenStudy (anonymous):

I mean, I know that but I dunno what to simplify. This? 2.5/2.95

hero (hero):

You said your answer choices were fractions. I don't know if you computed that or if you're just guessing from your list of answer choices.

OpenStudy (anonymous):

I computed that. Do you wanna see my choices?

hero (hero):

If you actually computed that, you need to show me your work

OpenStudy (anonymous):